# The universal multiplicity function: counting halos and voids

### v1

@@ Appendix C Global and boundary behavior
$\begin{split}&\det{\bf\Sigma}\simeq s(SD-1/4)=s\Gamma_{\delta\delta}\\ &\begin{array}[]{lcl}\Gamma_{BB}\simeq sD&&\Gamma_{B\delta^{\prime}}\simeq-s/2\\ \Gamma_{\delta^{\prime}\delta^{\prime}}\simeq sS&&\Gamma_{B\delta}\simeq-CD\\ \Gamma_{\delta\delta}=SD-1/4&&\Gamma_{\delta^{\prime}\delta}\simeq C/2\,,\end{array}\end{split}$ (C14)
@@
We start considering the first term on the right-hand side of Eq. (3), $I_{1}$, i.e. the second and third lines. The argument of the error function goes as
@@
$\rightarrow\frac{1}{\sqrt{2s}}\left(B(s)-\frac{DB_{S}-B^{\prime}_{S}/2}{\Gamma_{\delta\delta}}C\right).$ (C15)
The behavior at $s\rightarrow 0$ depends on the barrier shape. Taking into account a barrier of the form Eq. (16), in both the cases $B(s)\rightarrow{\rm const}$, i.e. $\gamma\geq 0$, and $B(s)\rightarrow\infty$, i.e. $\gamma<0$, as $s\rightarrow 0$, the above quantity goes to $\infty$, and the corresponding error function to 1. Using Eqs. (C14), the factor multiplying the exponential and the curly brackets behaves as
$\frac{\det{\bf\Sigma}}{4\pi\sqrt{\Gamma_{\delta\delta}}\Gamma_{\delta^{\prime}\delta^{\prime}}}\rightarrow\frac{\sqrt{\Gamma_{\delta\delta}}}{4\pi S},$ (C16)
and, therefore, since the exponential term does not depend on $s$, the first term of Eq. (3) at $s\rightarrow 0$ is
$I_{1}\rightarrow\frac{\sqrt{\Gamma_{\delta\delta}}}{2\pi S}\exp\left[-\frac{SB^{\prime 2}_{S}+DB_{S}^{2}-B^{\prime}_{S}B_{S}}{2\Gamma_{\delta\delta}}\right]$ (C17)
Let us consider the second term, $I_{2}$, i.e. the fourth line of Eq. (3). The factor multiplying the exponential behaves as
$\frac{\Gamma_{\delta^{\prime}\delta}}{4\pi S\sqrt{\Gamma_{\delta^{\prime}\delta^{\prime}}}}\rightarrow\frac{C}{8\pi S\sqrt{Ss}}\propto\sqrt{s},$ (C18)
@@
$SB^{2}(s)/2\Gamma_{\delta^{\prime}\delta^{\prime}}$ $\rightarrow B^{2}(s)/s\rightarrow\infty$
$sB_{S}^{2}/2\Gamma_{\delta^{\prime}\delta^{\prime}}$ $\rightarrow B_{S}^{2}/2S={\rm const}$ (C19)
@@
where the first equation goes always to $\infty$ for any behavior of $B(s)$, considering a form as Eq. (16), while the last equations may got either to a constant value or $\infty$ depending on $B(s)$ behavior. In any case, the exponent term tends to $\infty$- ∞ as $s\rightarrow 0$, so $I_{2}\rightarrow 0$.
The last term on the right-hand side of Eq. (3), i.e. $4^{\rm th}-5^{\rm th}$, is composed of a constant exponential term multiplying the sum $I_{3}+I_{4}$. The argument of the error function of $I_{4}$, i.e. $5^{\rm th}$ line of Eq. (3), in the limit $s\rightarrow 0$, behaves as
$\sqrt{\frac{S}{2\Gamma_{\delta^{\prime}\delta^{\prime}}}}\left(B(s)-\frac{C}{S}B_{S}\right)\rightarrow\frac{B(s)-CB_{S}/S}{\sqrt{2s}}\propto\frac{B(s)}{\sqrt{s}},$ (C20)
which goes to $\infty$ for any behavior of $B(s)$, if the functional form of Eq. (16) is considered. It follows that the corresponding error function goes to 1 and
$I_{4}\rightarrow\frac{B_{S}/2S-B^{\prime}_{S}}{2\sqrt{2\pi S}}$ (C21)
@@
$\frac{2\det\Sigma}{4\pi\sqrt{\Gamma_{\delta^{\prime}\delta^{\prime}}}\Gamma_{\delta^{\prime}\delta}}\rightarrow\frac{\Gamma_{\delta\delta}}{C}\sqrt{\frac{s}{S}}\rightarrow\frac{\Gamma_{\delta\delta}}{\sqrt{S}\partial_{s}C|_{s=0}}\frac{1}{\sqrt{s}},$ (C22)
multiplying ${\cal F}(x,\alpha,\beta)$. Let us consider the $x$, $\alpha$, $\beta$ quantities of Eqs. (14). The $x$ quantity behaves as
$x\rightarrow\frac{\partial_{s}C|_{s=0}}{2\sqrt{2S\Gamma_{\delta\delta}}}\left(B(s)-\frac{B_{S}}{\partial_{s}C}+\frac{2S}{\partial_{s}C}B^{\prime}_{S}\right),$ (C23)
@@
$\alpha=\frac{S\det{\bf\Sigma}}{\Gamma_{\delta^{\prime}\delta}^{2}}\rightarrow\frac{4\Gamma_{\delta\delta}Ss}{C^{2}}\rightarrow\frac{4\Gamma_{\delta\delta}S}{(\partial_{s}C|_{s=0})^{2}}\frac{1}{s}$ (C24)
@@
$\beta\rightarrow\sqrt{\frac{S}{2\Gamma_{\delta\delta}}}\left(\frac{B_{S}}{2S}-B^{\prime}_{s}\right)={\rm const}$ (C25)
as $s\rightarrow 0$. From the above quantities, it follows that the exponential $e^{-\alpha(t+\beta)^{2}}$ appearing in the integrand of ${\cal F}(x,\alpha,\beta)$, Eq. (13), behaves as an infinitesimal narrow Gaussian centered in $-\beta$. In the regime we are considering, it can be shown comparing $x$ and $\beta$, together with Eqs. (C14), that $-\beta<x$, and therefore
@@
$\rightarrow\frac{B_{S}/2S-B^{\prime}_{S}}{2\sqrt{2\pi S}}{\rm erf}\left[\sqrt{\frac{S}{2\Gamma_{\delta\delta}}}\left(\frac{B_{S}}{2S}-B^{\prime}_{s}\right)\right].$ (C26)
@@
Now it is possible to sum all the terms of Eq. (3) in the $s\rightarrow 0$ limits, from which it follows that $\lim_{s\rightarrow 0}{\cal P}(s)=\lim_{s\rightarrow S}{\cal P}(s)$ of Eq. (C10).
@@
Figure 8: Upper panel: ${\cal P}(s)/{\cal P}(S)$ for various $S$, listed in the legend. Bottom panel: ${\cal P}^{\prime}(s)/{\cal P}(s)$, the color refers to the same $S$ as in the upper panel.
Derivative of $\boldsymbol{{\cal P}(s)}$ and global behavior
@@
We now consider the global behavior of ${\cal P}(s)$. The upper panel of Fig. 8, shows the behavior of ${\cal P}(s)$ with respect to its value at the crossing scale, ${\cal P}(S)$, for various crossing scales $S$ corresponding to different color. The ${\cal P}(s)$ function shown here is the same as that described in Sec. 3, at $z=0$. It can be seen that $\lim_{s\rightarrow 0}{\cal P}(s)=\lim_{s\rightarrow S}{\cal P}(s)$, as previously derived, and that ${\cal P}(s)$ slowly varies in the range $0-S$, from sub- to a few percent level, depending on the corresponding crossing scale $S$. The lower panel shows the derivative ${\cal P}^{\prime}(s)$ with respect to ${\cal P}(s)$, for various crossing scales $S$. The derivative value is always within a few percent level the corresponding ${\cal P}(s)$ function. This further confirms that ${\cal P}(s)$ is a slowly varying function. From these properties, it follows that the integral of ${\cal P}(s)$ over $s$ can be approximated as $\int_{0}^{S}{\cal P}(s){\rm d}s\simeq S\lim_{s\rightarrow 0}{\cal P}(s)=S\lim_{s\rightarrow S}{\cal P}(s)$. The accuracy of this approximation increase as the crossing scale $S$ decreases.
Appendix D Measuring the moving barrier
In this Section we describe how the MCMC to fit $\alpha$, $\beta$, and $\gamma$ described in Sec. 4 is performed. We proceed in various steps. First, we obtain the multiplicity functions of the VSF measured in both the DEMNUni ICs and HR-DEMNUni ICs, for each of the void formation thresholds considered, Tab. 1. Then we consider the minimum useful void radius, in order to avoid scales in which voids are noise dominated. This minimum radius depends simultaneously on the spatial resolution of the simulation ICs, the void finder grid size, and the threshold value. In particular, we found that the void radius cannot be smaller than $\sim 5$ times the voxel side of the void finder grid. We found that 1000 grid division per box side guarantees the best possible spatial resolution given the number of particles of DEMNUni simulations, for both the standard and HR ICs, see Sec. 4. This results in a minimum useful radius of $\sim 10~{}h^{-1}{\rm Mpc}$ for DEMNUni ICs and $\sim 2.5~{}h^{-1}{\rm Mpc}$ for HR-DEMNUni ICs. Beyond the effect of the grid, the minimum useful void radius can be evaluated as a function of the scale corresponding to the Poissonian noise associated to the number of particles in a void with a given underdensity threshold. The expected number of particles in a sphere with radius $R$ is $N(R)=4\pi n_{\rm p}R^{3}/3$, where $n_{\rm p}$ is the mean number density of particles in the simulation box. The associated Poissonian uncertainty is $\sqrt{N}$. The number of particles contained in a void with formation threshold $\delta(z_{\rm IC})$ and radius $R$ detected by the void finder is $\Delta N(R)=\delta(z_{\rm IC})4\pi n_{\rm p}R^{3}/3$. We can define the radius $R_{\rm P}(P)$ as the radius of the void containing a number of particles equal to $P$ times the Poissonian uncertainty associated with the sphere with radius $R_{\rm P}(P)$: $\Delta N[R_{\rm P}(P)]=P\sqrt{N[R_{\rm P}(P)]}$. We use this radius as an estimator of the Poissonian uncertainty associated to the detection of a void with threshold $\delta(z_{\rm IC})$ and radius $R_{\rm P}(P)$
@@ Appendix D Measuring the moving barrier
$R_{\rm P}(P)=\left(\frac{P}{\delta(z_{\rm IC})}\right)^{2/3}\left(\frac{3}{4\pi n_{\rm p}}\right)^{1/3},$ (D1)

### v2

@@ C.2 $s\rightarrow 0$ limit
$\begin{split}&\det{\bf\Sigma}\simeq s(SD-1/4)=s\Gamma_{\delta\delta}\\ &\begin{array}[]{lcl}\Gamma_{BB}\simeq sD&&\Gamma_{B\delta^{\prime}}\simeq-s/2\\ \Gamma_{\delta^{\prime}\delta^{\prime}}\simeq sS&&\Gamma_{B\delta}\simeq-CD\\ \Gamma_{\delta\delta}=SD-1/4&&\Gamma_{\delta^{\prime}\delta}\simeq C/2\,,\end{array}\end{split}$ (C.14)
@@
We start considering the first term on the right-hand side of Eq. (3.2), $I_{1}$, i.e. the second and third lines. The argument of the error function goes as
@@
$\rightarrow\frac{1}{\sqrt{2s}}\left(B(s)-\frac{DB_{S}-B^{\prime}_{S}/2}{\Gamma_{\delta\delta}}C\right).$ (C.15)
The behaviour at $s\rightarrow 0$ depends on the barrier shape. Taking into account a barrier of the form Eq. (4.1), in both the cases $B(s)\rightarrow{\rm const}$, i.e. $\gamma\geq 0$, and $B(s)\rightarrow\infty$, i.e. $\gamma<0$, as $s\rightarrow 0$, the above quantity goes to $\infty$, and the corresponding error function to 1. Using Eqs. (C.14), the factor multiplying the exponential and the curly brackets behaves as
$\frac{\det{\bf\Sigma}}{4\pi\sqrt{\Gamma_{\delta\delta}}\Gamma_{\delta^{\prime}\delta^{\prime}}}\rightarrow\frac{\sqrt{\Gamma_{\delta\delta}}}{4\pi S},$ (C.16)
and, therefore, since the exponential term does not depend on $s$, the first term of Eq. (3.2) at $s\rightarrow 0$ is
$I_{1}\rightarrow\frac{\sqrt{\Gamma_{\delta\delta}}}{2\pi S}\exp\left[-\frac{SB^{\prime 2}_{S}+DB_{S}^{2}-B^{\prime}_{S}B_{S}}{2\Gamma_{\delta\delta}}\right]$ (C.17)
Let us consider the second term, $I_{2}$, i.e. the fourth line of Eq. (3.2). The factor multiplying the exponential behaves as
$\frac{\Gamma_{\delta^{\prime}\delta}}{4\pi S\sqrt{\Gamma_{\delta^{\prime}\delta^{\prime}}}}\rightarrow\frac{C}{8\pi S\sqrt{Ss}}\propto\sqrt{s},$ (C.18)
@@
$SB^{2}(s)/(2\Gamma_{\delta^{\prime}\delta^{\prime}})$ $\rightarrow B^{2}(s)/s\rightarrow\infty$
$sB_{S}^{2}/(2\Gamma_{\delta^{\prime}\delta^{\prime}})$ $\rightarrow B_{S}^{2}/2S={\rm const}$ (C.19)
@@
where the first equation goes always to $\infty$ for any behaviour of $B(s)$, considering a form as Eq. (4.1), while the last equations may got either to a constant value or $\infty$ depending on $B(s)$ behaviour. In any case, the exponent term tends to $\infty$- ∞ as $s\rightarrow 0$, so $I_{2}\rightarrow 0$.
The last term on the right-hand side of Eq. (3.2), i.e. $4^{\rm th}-5^{\rm th}$ lines of Eq. (3.2), is composed of a constant exponential term multiplying the sum $I_{3}+I_{4}$. The argument of the error function of $I_{4}$, i.e. $5^{\rm th}$ line of Eq. (3.2), in the limit $s\rightarrow 0$, behaves as
$\sqrt{\frac{S}{2\Gamma_{\delta^{\prime}\delta^{\prime}}}}\left(B(s)-\frac{C}{S}B_{S}\right)\rightarrow\frac{B(s)-CB_{S}/S}{\sqrt{2s}}\propto\frac{B(s)}{\sqrt{s}},$ (C.20)
which goes to $\infty$ for any behaviour of $B(s)$, if the functional form of Eq. (4.1) is considered. It follows that the corresponding error function goes to 1 and
$I_{4}\rightarrow\frac{B_{S}/2S-B^{\prime}_{S}}{2\sqrt{2\pi S}}$ (C.21)
@@
$\frac{2\det\Sigma}{4\pi\sqrt{\Gamma_{\delta^{\prime}\delta^{\prime}}}\Gamma_{\delta^{\prime}\delta}}\rightarrow\frac{\Gamma_{\delta\delta}}{C}\sqrt{\frac{s}{S}}\rightarrow\frac{\Gamma_{\delta\delta}}{\sqrt{S}\partial_{s}C|_{s=0}}\frac{1}{\sqrt{s}},$ (C.22)
multiplying ${\cal F}(x,\alpha,\beta)$. Let us consider the $x$, $\alpha$, $\beta$ quantities of Eqs. (3.2). The $x$ quantity behaves as
$x\rightarrow\frac{\partial_{s}C|_{s=0}}{2\sqrt{2S\Gamma_{\delta\delta}}}\left(B(s)-\frac{B_{S}}{\partial_{s}C}+\frac{2S}{\partial_{s}C}B^{\prime}_{S}\right),$ (C.23)
@@
$\alpha=\frac{S\det{\bf\Sigma}}{\Gamma_{\delta^{\prime}\delta}^{2}}\rightarrow\frac{4\Gamma_{\delta\delta}Ss}{C^{2}}\rightarrow\frac{4\Gamma_{\delta\delta}S}{(\partial_{s}C|_{s=0})^{2}}\frac{1}{s}$ (C.24)
@@
$\beta\rightarrow\sqrt{\frac{S}{2\Gamma_{\delta\delta}}}\left(\frac{B_{S}}{2S}-B^{\prime}_{s}\right)={\rm const}$ (C.25)
as $s\rightarrow 0$. From the above quantities, it follows that the exponential $e^{-\alpha(t+\beta)^{2}}$ appearing in the integrand of ${\cal F}(x,\alpha,\beta)$, Eq. (3.12), behaves as an infinitesimal narrow Gaussian centered in $-\beta$. In the regime we are considering, it can be shown comparing $x$ and $\beta$, together with Eqs. (C.14), that $-\beta<x$, and therefore
@@
$\rightarrow\frac{B_{S}/2S-B^{\prime}_{S}}{2\sqrt{2\pi S}}{\rm erf}\left[\sqrt{\frac{S}{2\Gamma_{\delta\delta}}}\left(\frac{B_{S}}{2S}-B^{\prime}_{s}\right)\right].$ (C.26)
@@
Now it is possible to sum all the terms of Eq. (3.2) in the $s\rightarrow 0$ limits, from which it follows that $\lim_{s\rightarrow 0}{\cal P}(s)=\lim_{s\rightarrow S}{\cal P}(s)$, as appearing in Eq. (C.10).
@@
Figure 8: Upper panel: ${\cal P}(s)/{\cal P}(S)$ for various $S$, listed in the legend. Bottom panel: ${\cal P}^{\prime}(s)/{\cal P}(s)$, the colour refers to the same $S$ as in the upper panel.
C.3 Derivative of ${\cal P}(s)$ and global behaviour
@@
We now consider the global behaviour of ${\cal P}(s)$. The upper panel of Fig. 8, shows the behaviour of ${\cal P}(s)$ with respect to its value at the crossing scale, ${\cal P}(S)$, for various crossing scales $S$ corresponding to different colour. The ${\cal P}(s)$ function shown here is the same as that described in Sec. 3, at $z=0$. It can be seen that $\lim_{s\rightarrow 0}{\cal P}(s)=\lim_{s\rightarrow S}{\cal P}(s)$, as previously derived, and that ${\cal P}(s)$ slowly varies in the range $0-S$, from sub- to a few percent level, depending on the corresponding crossing scale $S$. The lower panel shows the derivative ${\cal P}^{\prime}(s)$ with respect to ${\cal P}(s)$, for various crossing scales $S$. The derivative value is always within a few percent level the corresponding ${\cal P}(s)$ function. This further confirms that ${\cal P}(s)$ is a slowly varying function. From these properties, it follows that the integral of ${\cal P}(s)$ over $s$ can be approximated as $\int_{0}^{S}{\cal P}(s){\rm d}s\simeq S\lim_{s\rightarrow 0}{\cal P}(s)=S\lim_{s\rightarrow S}{\cal P}(s)$ in the limit of small $S$, obtaining Eq. (3.2).
Appendix D Local bias expansion
For any multiplicity function with dependence on the formation threshold, the corresponding local bias expansion naturally arises in the peak-background split framework, as a response of the formation threshold on the background variation [95, 2, 96, 6, 97, 9]. This is the case for the effective barrier approach explored in this paper. A modification in the background density due to a long-wavelength perturbation can be treated as a modification of the formation threshold, $f(S,\delta|\Delta_{l})=f(S,\delta-\Delta|0)$, where $\Delta_{l}$ is the long-wavelength perturbation. Therefore, the local bias expansion can be written as
$b_{\rm N}(S)=\frac{(-1)^{N}}{f(S)}\frac{\partial^{N}f(S)}{\partial\delta^{N}}.$ (D.1)
Since the dependence with respect to the formation threshold is contained in the moving barrier $B_{S}$, we separate the derivative of $B_{S}$ with respect to $\delta$ and the derivative of $f(S)$ with respect to the barrier. From the above equations, the first bias terms are
$b_{1}(S)$ $=\frac{-1}{f(S)}\frac{\partial B_{S}}{\partial\delta}\frac{\partial}{\partial B_{S}}f(S)\,,$ (D.2)
$b_{2}(S)$ $=\frac{1}{f(S)}\left[\frac{\partial^{2}B_{S}}{\partial\delta^{2}}\frac{\partial}{\partial B_{S}}+\left(\frac{\partial B_{S}}{\partial\delta}\right)^{2}\frac{\partial^{2}}{\partial B_{S}^{2}}\right]f(S)\,,$
$b_{3}(S)$ $=\frac{-1}{f(S)}\Bigg{[}\frac{\partial^{3}B_{S}}{\partial\delta^{3}}\frac{\partial}{\partial B_{S}}+3\frac{\partial B_{S}}{\partial\delta}\frac{\partial^{2}B_{S}}{\partial\delta^{2}}\frac{\partial^{2}}{\partial B_{S}^{2}}+\left(\frac{\partial B_{S}}{\partial\delta}\right)^{3}\frac{\partial^{3}}{\partial B_{S}^{3}}\Bigg{]}f(S)\,.$
Note that the barrier is formally a function of two variables, $B_{S}=B(S,\delta)$. When implementing the above equations with the analytical multiplicity functions presented in this work, Eqs. (3.2)-(3.2), this has to be taken into account when computing the derivative of $B^{\prime}$ with respect to the barrier, which reads
$\frac{\partial B^{\prime}_{S}}{\partial B_{S}}$ $=\left.\frac{\partial B^{\prime}_{S}}{\partial B_{S}}\right|_{S}=\left[\frac{\partial B_{S}}{\partial\delta}\right]^{-1}\frac{\partial B^{\prime}_{S}}{\partial\delta},$ (D.3)
$\frac{\partial^{2}B^{\prime}_{S}}{\partial B_{S}^{2}}$ $=\left[\frac{\partial B_{S}}{\partial\delta}\right]^{-1}\frac{\partial^{2}B^{\prime}_{S}}{\partial\delta^{2}}-\left[\frac{\partial B_{S}}{\partial\delta}\right]^{-2}\frac{\partial^{2}B_{S}}{\partial\delta^{2}}\frac{\partial B^{\prime}_{S}}{\partial\delta},$
$\frac{\partial^{3}B^{\prime}_{S}}{\partial B_{S}^{3}}$ $=\left[\frac{\partial B_{S}}{\partial\delta}\right]^{-1}\frac{\partial^{3}B^{\prime}_{S}}{\partial\delta^{3}}-2\left[\frac{\partial B_{S}}{\partial\delta}\right]^{-2}\frac{\partial^{2}B_{S}}{\partial\delta^{2}}\frac{\partial^{2}B^{\prime}_{S}}{\partial\delta^{2}}+$
$\qquad\left\{2\left[\frac{\partial B_{S}}{\partial\delta}\right]^{-3}\left[\frac{\partial^{2}B_{S}}{\partial\delta^{2}}\right]^{2}-\left[\frac{\partial B_{S}}{\partial\delta}\right]^{-2}\frac{\partial^{3}B_{S}}{\partial\delta^{3}}\right\}\frac{\partial B^{\prime}_{S}}{\partial\delta}\,.$
When considering the approximated multiplicity function, Eq. (3.2), we obtain
@@ Appendix D Local bias expansion
$\frac{\partial f(S)}{\partial B_{S}}$ $=-\frac{B_{S}}{S}f(S)+\frac{e^{-B_{S}^{2}/2S}}{2\sqrt{2\pi S}}\partial_{B}\Delta_{B}\times\left[{\rm erf}\left(\sqrt{\frac{S}{2\Gamma_{\delta\delta}}}\Delta_{B}\right)+1\right],$
$\frac{\partial^{2}f(S)}{\partial B_{S}^{2}}$ $=\left(\frac{B_{S}^{2}}{S^{2}}-\frac{1}{S}\right)f(S)+\frac{e^{-B_{S}^{2}/2S}}{\sqrt{2\pi S}}\left\{(\partial_{B}\Delta_{B})^{2}\sqrt{\frac{S}{2\pi\Gamma_{\delta\delta}}}\exp\left(-\frac{S\Delta_{B}^{2}}{2\Gamma_{\delta\delta}}\right)+\right.$
$\left.\frac{1}{2}\left(\partial_{B}^{2}\Delta_{B}-2\frac{B}{S}\partial_{B}\Delta_{B}\right)\left[{\rm erf}\left(\sqrt{\frac{S}{2\Gamma_{\delta\delta}}}\Delta_{B}\right)+1\right]\right\},$ (D.4)
$\frac{\partial^{3}f(S)}{\partial B_{S}^{3}}$ $=\left(\frac{3B}{S^{2}}-\frac{B^{3}}{S^{3}}\right)f(S)+\frac{e^{-B_{S}^{2}/2S}}{\sqrt{2\pi S}}\times$
$\Bigg{\{}\left[3\partial^{2}_{B}\Delta_{B}\partial_{B}\Delta_{B}-\frac{3B}{S}(\partial_{B}\Delta_{B})^{2}-\frac{S}{\Gamma_{\delta\delta}}\Delta_{B}(\partial_{B}\Delta_{B})^{3}\right]\sqrt{\frac{S}{2\pi\Gamma_{\delta\delta}}}\exp\left(-\frac{S\Delta_{B}^{2}}{2\Gamma_{\delta\delta}}\right)+$
$\quad\frac{1}{2}\left[\left(\frac{3B}{S^{2}}-\frac{3}{S}\right)\partial_{B}\Delta_{B}-\frac{3B}{S}\partial^{2}_{B}\Delta_{B}+\partial^{3}_{B}\Delta_{B}\right]\left[{\rm erf}\left(\sqrt{\frac{S}{2\Gamma_{\delta\delta}}}\Delta_{B}\right)+1\right]\Bigg{\}}$
where $\Delta_{B}=(B_{S}/2S-B^{\prime}_{S})$, $\partial_{B}\Delta_{B}=1/2S-\partial B^{\prime}_{S}/\partial B_{S}$, $\partial^{2}_{B}\Delta^{2}_{B}=-\partial^{2}B^{\prime}_{S}/\partial B_{S}^{2}$, and $\partial^{3}_{B}\Delta^{3}_{B}=-\partial^{3}B^{\prime}_{S}/\partial B_{3}^{2}$.
Appendix E Measuring the moving barrier
[…]
$R_{\rm P}(P)=\left(\frac{P}{\delta(z_{\rm IC})}\right)^{2/3}\left(\frac{3}{4\pi n_{\rm p}}\right)^{1/3},$ (E.1)